OpenStudy (anonymous):

derive algebraically: e^-ln2

OpenStudy (anonymous):

\[e^{-\ln(2)}=e^{\ln(\frac{1}{2})}=\frac{1}{2}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw